# Prompt-Enhanced Spatio-Temporal Graph Transfer Learning

### v1

@@ Abstract.
Spatio-temporal graph neural networks have demonstrated efficacy in capturing complex dependencies for urban computing tasks such as forecasting and kriging. However, their performance is constrained by the reliance on extensive data for training on specific tasks, which limits their adaptability to new urban domains with varied demands. Although transfer learning has been proposed to address this problem by leveraging knowledge across domains, cross-task generalization remains underexplored in spatio-temporal graph transfer learning methods due to the absence of a unified framework. To bridge this gap, we propose Spatio-Temporal Graph Prompting (STGP), a prompt-enhanced transfer learning framework capable of adapting to diverse tasks in data-scarce domains. Specifically, we first unify different tasks into a single template and introduce a task-agnostic network architecture that aligns with this template. This approach enables the capture of spatio-temporal dependencies shared across tasks. Furthermore, we employ learnable prompts to achieve domain and task transfer in a two-stage prompting pipeline, enabling the prompts to effectively capture domain knowledge and task-specific properties at each stage. Extensive experiments demonstrate that STGP outperforms state-of-the-art baselines in three downstream tasks forecasting, kriging, and extrapolation by a notable margin.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Transfer learning, which focuses on transferring knowledge across domains and tasks, emerges as a promising machine learning approach in scenarios where training data is limited (Zhuang et al., 2020). It aims to leverage knowledge learned from a data-rich source domain to perform diverse downstream tasks in a data-scarce target domain. Recently, researchers have realized the data scarcity issue in spatio-temporal graphs and proposed a variety of solutions (Wang et al., 2019; Yao et al., 2019). As shown in Fig. 1(a)-(b), these models are first pre-trained on a source domain and then fine-tuned on a target domain. To achieve effective domain transfer, they are supported by a transfer model. For instance, CrossTReS (Jin et al., 2022) captures domain shift by a weighting mechanism that filters out exclusive patterns in the source domain. TPB (Liu et al., 2023e) stores shareable patterns across the domains by a memory bank. Despite the intriguing performance obtained by these transfer learning systems, they are specifically designed for the task of forecasting by incorporating task-specific architectures. Consequently, they fall short in achieving task transfer due to their sensitivity to changes in task specifications  (Brown et al., 2020). In this work, we aim to move towards more general task-agnostic frameworks that are capable of performing multiple tasks in the target domain.
@@
To tackle the problem of task transfer, an emerging technique known as prompt tuning has gained prominence (Brown et al., 2020). This concept, originating from Natural Language Processing (NLP), first unifies different tasks into a task template with a single training objective. Then, it leverages a Transformer model that is task-agnostic and employs prompts to instruct task-specific knowledge. Compared to model fine-tuning, prompt tuning gains more efficiency by freezing the pre-trained model. Meanwhile, it still yields performance that is comparable to, or even surpasses the fine-tuning methods (Liu et al., 2023d). Recently, the community has found that prompt tuning can also perform domain transfer by incorporating domain-specific prompts into the input data (Ge et al., 2023). Inspired by this paradigm, we argue that prompting can similarly be introduced for transfer learning in spatio-temporal graphs, primarily for two key reasons. First, prompting can suit different spatio-temporal tasks flexibly by leveraging a unified task template and task-agnostic architectures. Second, it provides an innovative idea to achieve both domain and task transfer efficiently from a data perspective: keeping the model frozen while adapting the data space of the target domain.
@@
However, applying prompt tuning to spatio-temporal graphs is non-trivial due to the following challenges. Firstly, to enable task transfer, the design of the model and the knowledge captured by it through pre-training should be task-agnostic. Taking language models (Min et al., 2023) as an example, they unify the language tasks into a template of word prediction and utilize Transformers as the learning model. The template and the network architecture ensure its applicability and the learning of a latent space shared by various tasks (Qin et al., 2021). Nevertheless, how do we design a model and a task template for spatio-temporal graphs that preserves shareable dependencies among tasks is under-explored. Secondly, existing prompting methods treat domain transfer and task transfer as distinct problems, introducing different solutions to address them by using either natural language or learnable prompts (Ge et al., 2023; Sun et al., 2023). However, spatio-temporal graphs lack an effective way to incorporate language that can prompt domain and task knowledge. Moreover, it is less apparent what mechanism the learnable prompts should take to guide both transfer objectives. Thus, how to learn prompts that capture both domain knowledge and task-specific properties is also a challenge waiting to be solved.
@@
To address the aforementioned challenges, we propose a novel Spatio-Temporal Graph Prompting framework, termed STGP, with the goal of facilitating domain transfer while enabling capabilities for multiple downstream tasks. First, we unify spatio-temporal graph tasks into a single task template, inspired by STEP (Shao et al., 2022), and introduce a task-agnostic network architecture that aligns with this template. Moreover, we introduce a two-stage prompting approach. As depicted in Fig. 1(c), the prompting method sequentially learns domain transfer and task transfer in its respective prompting stages.
In particular, to address the first challenge, we unify spatio-temporal graph tasks into a data reconstruction template, where each task is represented as predictions of data at specific masked positions in a spatio-temporal graph. To align with the template, we introduce a task-agnostic network, empowered by a Transformer-based encoder and a gated spatio-temporal decoder, that takes in input data to predict masked signals. In this work, we adopt the predictions of randomly masked data as the model’s pre-training task to facilitate the learning of shareable spatio-temporal dependencies. Then, as illustrated in Fig. 2, we apply the pre-trained model to two common spatio-temporal downstream tasks—forecasting and kriging—and introduce a new task—extrapolation. The extrapolation task aims at predicting future signals for nodes lacking historical data, which holds significant practical importance. For instance, it enables the use of limited sensors to forecast future traffic conditions across an entire city.
Regarding the second challenge, we decouple the transferring of domain and task knowledge and propose a prompting pipeline consisting of two stages. The initial stage focuses on domain transfer, dynamically adapting the target domain’s data space to the source domain used for model pre-training. In practice, we freeze the pre-trained model, forward the domain-prompted data into the encoder, and train the prompts using the same task during pre-training. In the second stage, domain prompts are fixed and task prompts operate on the representations produced by the encoder. The task-prompted representations are fed into the decoder, which enables it to differentiate various downstream tasks and reach optimal performance for each task. Our model is more efficient than current spatio-temporal transfer learning methods by leveraging domain prompts to transfer knowledge without the need for fine-tuning the pre-trained model. Meanwhile, by employing task prompting, our STGP shows versatility across multiple downstream tasks, yielding superior performance compared to other baselines. To summarize, our contributions lie in three aspects:
@@
  • We introduce a comprehensive framework for spatio-temporal graph transfer learning called STGP. STGP leverages a unified template and a task-agnostic architecture that transfers the model to multiple downstream tasks in the target domain, such as forecasting, kriging, and the new task of extrapolation.
@@
  • We perform extensive experiments across four public datasets. The results and discussions highlight the superior performance of STGP and the efficacy of each component of the model.
@@ 2.1. Spatio-Temporal Graph Neural Networks
STGNNs have become one of the dominant paradigms for urban computing (Jin et al., 2023b), leading to a series of applications, such as forecasting and kriging. DCRNN (Li et al., 2018) and STGCN (Yu et al., 2018) are two pioneer methods for forecasting, which employ GNNs to capture spatial relations and deploy sequential models to model temporal dependencies. Graph WaveNet (Wu et al., 2019) proposes a matrix learning module to model hidden structural information, which becomes an influential idea for other methods (Lan et al., 2022; Shao et al., 2022; Rao et al., 2022). In terms of the task of kriging, a growing number of models have been recently developed to address it. For instance, KCN (Appleby et al., 2020) applies GNNs to capture dependencies between existing and unknown nodes, while IGNNK (Wu et al., 2021a) further considers temporal relations. DualSTN (Hu et al., 2023) learns joint correlations between time and space. SATCN (Wu et al., 2021b) and INCREASE (Zheng et al., 2023) propose to capture statistical graph information. However, STGNNs heavily rely on sufficient training data and are designed exclusively for a specific task. In this study, we are devoted to addressing these challenges through the proposed prompt-based framework.
@@ 2.2. Spatio-Temporal Transfer learning
To improve the performance of STGNNs in scenarios with limited data, transfer learning methods are employed for spatio-temporal data. The model undergoes initial pre-training on a data-rich source domain and is then transferred to a target domain where data is scarce. Additionally, a transfer module is introduced to facilitate domain adaptation. RegionTrans (Wang et al., 2019), MetaST (Pan et al., 2019), and CrossTReS (Jin et al., 2022) focus on grid data, learning region correlations between target and source cities. In the graph setting, ST-GFSL (Lu et al., 2022) introduces to generate network parameters specialized for the target domain. TPB (Liu et al., 2023e) learns temporal patterns shareable across the domains. TransGTR (Jin et al., 2023a) proposes to generate graph structures for the target domain that exhibit structural distributions akin to the source domain. Although these approaches introduce various transfer mechanisms to enhance model performance, most of them are optimized through a complex meta-learning framework such as Reptile (Nichol et al., 2018) or MAML (Finn et al., 2017) that require fine-tuning of the pre-trained model, significantly diminishing their efficiency. Moreover, their exclusive focus on forecasting, using task-specific architectures, results in a lack of versatility for multiple downstream tasks.
@@ 2.3. Prompt Tuning
Originating from Natural Language Processing, prompt tuning methods have gained widespread use for adapting pre-trained language models to a diverse range of downstream tasks (Liu et al., 2023d). These methods formulate the pre-training and downstream tasks into a single template of word prediction and employ learnable or natural language prompts to instruct task-related properties. Recently, researchers have discovered that prompting shows significant potential in domain transfer by guiding the model with domain knowledge (Ge et al., 2023). Aside from the language tasks, the prompting methods have also gained attention from various research areas such as Computer Vision (Jia et al., 2022), Graph Learning (Sun et al., 2023), and Time Series Modeling (Liu et al., 2023a; Xue and Salim, 2023). Despite their success, prompt-based methods remain under-explored in the context of spatio-temporal graphs, with few methods proposed (Zhang et al., 2023; Li et al., 2024). Furthermore, the transfer learning scenario demands simultaneous management of domain transfer and multi-task capabilities within a unified framework, thereby amplifying the complexity of the challenge.
@@ 4. Methodology
In this paper, we aim to introduce a spatio-temporal graph transfer learning framework, empowered by prompt tuning, that can be applied to multiple tasks. To achieve this goal, we first consolidate the pre-training and downstream tasks into a unified template as illustrated in Fig. 3(c). Next, we design a network that aligns with the template (Fig. 3(a)) and propose a two-stage prompting pipeline for domain and task transfer (Fig. 3(b)), respectively. In the following sections, we detail each component of STGP.
@@ 4.1. Unification Spatio-Temporal Graph Tasks
The effectiveness of applying a pre-trained model to diverse downstream tasks greatly relies on a unified task template, which ensures the learning of a shareable latent space across tasks. However, it remains unclear how to unify these tasks for spatio-temporal graphs. As suggested by (Jin et al., 2023c), different tasks place differing emphases on spatial or temporal dimensions. For instance, forecasting predominantly relies on the temporal evolution of historical signals, while kriging focuses more on the spatial relationships between observed and unobserved nodes. Consequently, existing models have to adopt task-specific training objectives and architectures to align with these properties (Wu et al., 2021b; Lu et al., 2022), which constrains their applicability on multiple downstream tasks.
Inspired by STEP (Shao et al., 2022) and GraphMAE (Hou et al., 2022), we reformulate the pre-training and downstream tasks as the reconstruction of masked data. As illustrated in Fig. 3(c), this unified template involves predicting masked data and requires the extraction of useful spatio-temporal dependencies from unmasked signals. In this way, these tasks can be viewed as unique, each characterized by masks placed at specific positions. For example, forecasting predicts masked signals at future time steps, while kriging aims to infer masked data that represents unobserved nodes. We adopt the predictions of randomly masked data as the pre-training task, which will be detailed below. Since the data masking relies solely on the task definition, without prioritizing spatial or temporal factors, our unified template ensures equal emphasis on dependencies across both dimensions, leading to consistent performance across all tasks.
@@ 4.2. Pre-Training and Spatio-Temporal Networks
Our pre-training approach involves predicting randomly masked data, propelling the model to identify shareable spatio-temporal dependencies for downstream tasks. To facilitate this, we introduce a spatio-temporal graph network featuring an encoder-decoder architecture: the Transformer-based encoder captures complex dependencies within unmasked data, while the gated decoder focuses on the predictions of masked data.

### v2

# Prompt-Based Spatio-Temporal Graph Transfer Learning

@@ Abstract.
Spatio-temporal graph neural networks have proven efficacy in capturing complex dependencies for urban computing tasks such as forecasting and kriging. Yet, their performance is constrained by the reliance on extensive data for training on a specific task, thereby limiting their adaptability to new urban domains with varied task demands. Although transfer learning has been proposed to remedy this problem by leveraging knowledge across domains, the cross-task generalization still remains under-explored in spatio-temporal graph transfer learning due to the lack of a unified framework. To bridge the gap, we propose Spatio-Temporal Graph Prompting (STGP), a prompt-based framework capable of adapting to multi-diverse tasks in a data-scarce domain. Specifically, we first unify different tasks into a single template and introduce a task-agnostic network architecture that aligns with this template. This approach enables capturing dependencies shared across tasks. Furthermore, we employ learnable prompts to achieve domain and task transfer in a two-stage prompting pipeline, facilitating the prompts to effectively capture domain knowledge and task-specific properties. Our extensive experiments demonstrate that STGP outperforms state-of-the-art baselines in three tasks—forecasting, kriging, and extrapolation—achieving an improvement of up to 10.7%.
Transfer learning; prompt tuning; spatio-temporal data mining
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679554isbn: 979-8-4007-0436-9/24/10ccs: Information systems Spatial-temporal systems
@@ 1. Introduction
Transfer learning, which focuses on transferring knowledge across domains and tasks, emerges as a promising machine learning approach in scenarios where training data is limited (Zhuang et al., 2020). It aims to leverage knowledge learned from a data-rich source domain to perform diverse downstream tasks in a data-scarce target domain. Nowadays, researchers have realized the data scarcity issue in spatio-temporal graphs and proposed a variety of solutions (Wang et al., 2019; Yao et al., 2019). As shown in Fig. 1(a)-(b), these models are first pre-trained on a source domain and then fine-tuned on a target domain. To achieve effective domain transfer, they are further supported by a transfer module. For instance, CrossTReS (Jin et al., 2022) captures domain shift by a weighting mechanism that filters out patterns exclusively in the source domain. TPB (Liu et al., 2023e) stores shareable patterns across the domains by a memory bank. Despite the intriguing performance obtained by these systems, they are specifically designed for the task of forecasting by incorporating task-specific architectures. Consequently, they fall short in achieving task transfer due to their sensitivity to changes in task specifications  (Brown et al., 2020). In this work, we aim to move towards general task-agnostic frameworks that are capable of performing multiple tasks in the target domain.
@@
To tackle the problem of task transfer, an emerging technique known as prompt tuning has gained prominence (Brown et al., 2020). This concept, originating from Natural Language Processing (NLP), first unifies different tasks into one task template with a single training objective. Then, it leverages pre-trained Transformers that are task-agnostic and employs prompts to instruct task-specific knowledge. Compared to fine-tuning, prompt tuning gains more efficiency by freezing the pre-trained model. Meanwhile, it still yields performance that is comparable to, or even surpasses the fine-tuning methods (Liu et al., 2023d). Recently, the community has found that prompt tuning can also perform domain transfer by incorporating domain-specific prompts into the input data (Ge et al., 2023). Inspired by this paradigm, we argue that prompting can similarly be introduced for transfer learning in spatio-temporal graphs, primarily for two key reasons. First, prompting can suit different spatio-temporal tasks flexibly by leveraging a unified task template and task-agnostic architectures. Second, it provides an innovative idea to achieve both domain and task transfer efficiently from a data perspective: keeping the model frozen while adapting the data space of the target domain.
@@
However, applying prompt tuning to spatio-temporal graphs is non-trivial due to the following challenges. Firstly, to enable task transfer, the design of the model and the knowledge captured by it through pre-training should be task-agnostic. Taking language models (Min et al., 2023) as an example, they unify the language tasks into a template of word prediction and utilize Transformers as the learning model. The template and the network architecture ensure its applicability and the learning of latent space shared by various tasks (Qin et al., 2021). Nevertheless, how do we design a model and a task template for spatio-temporal graphs that preserves shareable dependencies among tasks is under-explored. Secondly, existing prompting methods treat domain transfer and task transfer as distinct problems, introducing different solutions to address them by using either natural language or learnable prompts (Ge et al., 2023; Sun et al., 2023). However, spatio-temporal graphs lack an effective way to incorporate language that can prompt domain and task knowledge. Moreover, it is less apparent what mechanism the learnable prompts should take to guide both transfer objectives. Thus, how to learn prompts that capture both domain knowledge and task-specific properties is also a challenge waiting to be solved.
@@
To address the aforementioned challenges, we propose a novel Spatio-Temporal Graph Prompting framework, termed STGP, with the goal of facilitating domain transfer while enabling capabilities for multiple downstream tasks. First, we unify spatio-temporal graph tasks into a single task template, inspired by STEP (Shao et al., 2022), and introduce a task-agnostic network that aligns with this template. Moreover, we introduce a two-stage prompting approach. As depicted in Fig. 1(c), the prompting method sequentially learns domain transfer and task transfer in its respective prompting stages.
@@
In particular, to address the first challenge, we unify spatio-temporal graph tasks into a data reconstruction template, where each task is represented as predictions of data at specific masked spatio-temporal positions. To align with the template, we introduce a task-agnostic network, empowered by a Transformer-based encoder and a gated spatio-temporal decoder, that takes in input data to predict masked signals. In this work, we adopt predictions of randomly masked data as the model’s pre-training task to facilitate the learning of shareable spatio-temporal dependencies. Then, as illustrated in Fig. 2, we apply the pre-trained model to two common downstream tasks—forecasting and kriging—and introduce a new task—extrapolation. The extrapolation task aims at predicting future signals for nodes lacking historical data, which holds significant practical importance. For example, it enables the use of limited sensors to forecast future traffic conditions across an entire city.
Regarding the second challenge, we decouple the transfer of domain and task knowledge and propose a prompting pipeline consisting of two stages. The initial stage focuses on domain transfer, adapting the target domain’s data space to the source domain used for model pre-training. In practice, we freeze the pre-trained model, forward the domain-prompted data into the encoder, and train the prompts using the pre-training task. In the second stage, domain prompts are fixed and task prompts operate on the representations produced by the encoder. The task-prompted representations are fed into the decoder, enabling it to differentiate downstream tasks and reach optimal performance for each task. Our model is more efficient than current spatio-temporal transfer learning methods by leveraging domain prompts to transfer knowledge without the need for fine-tuning the pre-trained model. Meanwhile, by employing task prompting, our STGP shows versatility across multiple downstream tasks, yielding superior performance compared to other baselines. To summarize, our contributions lie in three aspects:
  • We introduce STGP, a novel framework for spatio-temporal graph transfer learning that leverages a unified template and a task-agnostic network to transfer knowledge across multiple tasks in the target domain, such as forecasting, kriging, and extrapolation.
@@
  • We conduct extensive experiments across four public datasets. The results and discussions highlight the superior performance of STGP and the efficacy of each component of the model.
@@ 2.1. Spatio-Temporal Graph Neural Networks
STGNNs have become one of the dominant paradigms for urban computing (Jin et al., 2023b), leading to a series of applications, such as forecasting and kriging. DCRNN (Li et al., 2018) and STGCN (Yu et al., 2018) are two pioneer methods for forecasting, which employ GNNs to capture spatial relations and deploy sequential models to model temporal dependencies. Graph WaveNet (Wu et al., 2019) proposes a matrix learning module to model hidden structural information, which becomes an influential idea for following methods (Lan et al., 2022; Shao et al., 2022; Rao et al., 2022). In terms of the task of kriging, a growing number of models have been recently developed to address it. For instance, KCN (Appleby et al., 2020) applies GNNs to capture dependencies between existing and unknown nodes, while IGNNK (Wu et al., 2021a) further considers temporal relations. DualSTN (Hu et al., 2023) learns joint correlations between time and space. SATCN (Wu et al., 2021b) and INCREASE (Zheng et al., 2023) propose to capture statistical graph information. However, STGNNs heavily rely on sufficient training data and are designed exclusively for a specific task. In this study, we are devoted to addressing these challenges through the proposed prompt-based framework.
@@ 2.2. Spatio-Temporal Transfer learning
To improve the performance of STGNNs in scenarios with data scarcity, transfer learning methods are employed for spatio-temporal data. The model undergoes initial pre-training on a data-rich source domain and is then transferred to a target domain where data is scarce. Additionally, a transfer module is introduced to facilitate domain adaptation. RegionTrans (Wang et al., 2019), MetaST (Pan et al., 2019), and CrossTReS (Jin et al., 2022) focus on grid data, learning region correlations between target and source cities. In the graph setting, ST-GFSL (Lu et al., 2022) introduces to generate network parameters specialized for the target domain. TPB (Liu et al., 2023e) learns temporal patterns shareable across the domains. TransGTR (Jin et al., 2023a) proposes to generate graph structures for the target domain that exhibit structural distributions akin to the source domain. Although these approaches introduce various transfer mechanisms to enhance model performance, most of them are optimized through a complex meta-learning framework such as Reptile (Nichol et al., 2018) or MAML (Finn et al., 2017) that require fine-tuning of the pre-trained model, which significantly diminishes their efficiency. Moreover, their exclusive focus on forecasting, using task-specific architectures, results in a lack of versatility for multiple downstream tasks.
@@ 2.3. Prompt Tuning
Originating from Natural Language Processing, prompt tuning methods have gained widespread use for adapting pre-trained language models to a range of downstream tasks (Liu et al., 2023d). These methods formulate the pre-training and downstream tasks into a single template of word prediction and employ learnable or natural language prompts to instruct task-related properties. In recent times, researchers have discovered that prompting shows potential in domain transfer by guiding the model with domain knowledge (Ge et al., 2023). Aside from the language tasks, prompting methods have also gained attention from various research areas such as Computer Vision (Jia et al., 2022), Graph Learning (Sun et al., 2023), and Time Series Modeling (Liu et al., 2023a; Xue and Salim, 2023). Despite their success, prompt-based methods remain under-explored in the context of spatio-temporal learning, with few proposed methods for graph (Li et al., 2024) and grid data (Yuan et al., 2024) that focus on domain transfer only. Nevertheless, the transfer learning scenario demands simultaneous management of domain transfer and multi-task capabilities, thereby amplifying the complexity of the challenge.
@@ 4. Methodology
In this paper, we aim to introduce a spatio-temporal graph transfer learning framework, empowered by prompt tuning, that can be applied to multiple tasks. To achieve this goal, we first consolidate the pre-training and downstream tasks into a unified template as illustrated in Fig. 3(c). Next, we design a network architecture that aligns with the template (Fig. 3(a)) and propose a two-stage prompting pipeline for domain and task transfer (Fig. 3(b)), respectively. In the following sections, we detail each component of STGP.
@@ 4.1. Unification Spatio-Temporal Graph Tasks
The effectiveness of applying a pre-trained model to diverse downstream tasks greatly relies on a unified task template, which ensures the learning of shareable latent space across tasks. However, it remains unclear how to unify these tasks for spatio-temporal graphs. As suggested by (Jin et al., 2023c), different tasks place differing emphases on spatial and temporal dimensions. For instance, forecasting predominantly relies on the temporal evolution of historical signals, while kriging focuses more on the spatial relationships between observed and unobserved nodes. Consequently, existing models have to adopt task-specific training objectives and architectures to align with these properties (Wu et al., 2021b; Lu et al., 2022), which constrains their applicability on multiple downstream tasks.
Inspired by STEP (Shao et al., 2022) and GraphMAE (Hou et al., 2022), we reformulate the pre-training and downstream tasks as the reconstruction of masked data. As illustrated in Fig. 3(c), this unified template involves predicting masked data that requires the extraction of useful spatio-temporal dependencies from unmasked signals. In this way, these tasks can be defined when they are characterized by masks placed at specific positions. For example, forecasting predicts masked signals at future time steps, while kriging aims to infer masked data that represents unobserved nodes. We adopt the predictions of randomly masked data as the pre-training task, which will be detailed below. Since the data masking relies solely on the task definition, without prioritizing spatial or temporal factors, our unified template ensures equal emphasis on dependencies across both dimensions, leading to consistent performance across all tasks.
@@ 4.2. Pre-Training and Spatio-Temporal Networks
Our pre-training approach involves predicting randomly masked data, propelling the model to identify shareable spatio-temporal dependencies for downstream tasks. To facilitate this, we introduce a spatio-temporal graph network featuring an autoencoder that comprises a Transformer-based encoder to capture complex dependencies within unmasked data and a gated decoder focusing on the predictions of masked data.